### v2

# Human-Agent Joint Learning for Efficient Robot Manipulation Skill Acquisition

Abstract
Employing a teleoperation system for gathering demonstrations offers the potential for more efficient learning of robot manipulation. However, teleoperating a robot arm equipped with a dexterous hand or gripper, via a teleoperation system poses significant challenges due to its high dimensionality, complex motions, and differences in physiological structure. In this study, we introduce a novel system for joint learning between human operators and robots, that enables human operators to share control of a robot end-effector with a learned assistive agent, facilitating simultaneous human demonstration collection and robot manipulation teaching. In this setup, as data accumulates, the assistive agent gradually learns. Consequently, less human effort and attention are required, enhancing the efficiency of the data collection process. It also allows the human operator to adjust the control ratio to achieve a trade-off between manual and automated control. We conducted experiments in both simulated environments and physical real-world settings. Through user studies and quantitative evaluations, it is evident that the proposed system could enhance data collection efficiency and reduce the need for human adaptation while ensuring the collected data is of sufficient quality for downstream tasks. Videos are available at https://norweig1an.github.io/human-agent-joint-learning.github.io/.
Keywords: Robotic Teleoperation, Robot Manipulation, Imitation Learning
footnotetext: * denotes equal contribution, †denotes corresponding author.
1 Introduction
The long-term vision in the field of robot learning has been to enable robots to perform diverse tasks at a human level in the physical world. Recently, significant progress has been made toward this goal by learning robot manipulation policies from demonstrations. Previous studies have utilized teleoperation systems [1, 2, 3, 4, 5, 6] to collect human demonstrations, and learning-based policies [7, 8, 9] have been formulated using the gathered data. Despite the notable advancements, several challenges still need to be addressed. For example, in vision-based teleoperation systems, even with state-of-the-art 3D hand pose estimation algorithms [10, 11, 12, 13], errors persist that significantly affect the teleoperation. Additionally, discrepancies between the structures of human hands and robot end-effectors, along with the lack of haptic feedback during contact-rich manipulation, also pose challenges. As a result, current teleoperation systems require human operators to practice extensively to adapt to these differences and gather the necessary data. This means that many human adaptations are essential. Furthermore, to meet the data requirements of robotic systems, humans need to collect large amounts of data, making this process very burdensome.
Figure 1: Traditional frameworks typically separate human and agent training, requiring operators first to learn the task environment before data collection. This often leads to inefficiencies due to delayed and insufficient data gathering. In our framework, we integrate human and agent training from the start in a joint learning model. This enables simultaneous development and adapts the agents to human operation more effectively, enhancing overall efficiency and promoting better collaboration between humans and machines allowing for human effortless adaptation data collection.
Naturally, a question was raised: in data-collection, how to make human adaptation less or even free while keeping the data quality? Here, we aim to address this question and argue that human-agent joint learning can help. That said, an effective and efficient teleoperation system should be designed to preferentially capture the operator’s intentions for directing a robot end effector and pose the main frame, while concurrently enabling an autonomous agent to help us ensure motion stability and interpolate the details. To this end, we propose a framework that achieves shared control between the human and a learned assistive agent. As shown in Fig. 1, our human effortless adaptation framework seeks to balance human needs and system performance, aiming to directly enhance the efficiency and quality of data collection by reducing the time and effort in human adaptation.
Given our human-agent joint learning approach, we allow the data acquisition agent to grow and learn along with the human operator, reducing the learning burden of the human operator throughout the data collection task. Inspired by shared autonomy [14, 15, 16], we introduce a novel teleoperation system that enables collaboration between humans and learning-based agents to control a robot jointly. In particular, our proposed system provides the flexibility to adjust a “control ratio” between the human operator and a learning-based agent. A lower control ratio, in the beginning, signifies a greater emphasis on humans to teach the agent the finer-grained knowledge under the structure of human intention and principal actions. As the agent’s learning improves, a higher ratio later indicates increased autonomy from the learned agent to replace the human effort to “inpaint” the whole process given only human intention and principal actions. Additionally, once a sufficient amount of data is gathered, we have the option to transition the shared autonomy agent to full autonomy by reducing the human control ratio to zero.
We implement our framework as a joint learning system, which facilitates a human operator’s intention and an assistive agent’s execution of teleoperation tasks. For human operators, their inputs consist of intuitive control actions based on visual feedback and past cognitive experience. For the agent, their inputs are derived from a combination of sensors, data streams, and possibly pre-processed information. This can include visual data from cameras [1], tactile data from force sensors [3], and any additional context provided by the system [6], such as object recognition or environmental mapping. The agent processes these inputs using algorithms designed to interpret the task requirements and generate actions to assist the human operator, utilizing a diffusion model [8, 17, 18, 19] as the backbone enables us to adjust the control ratio by modifying the step number of the forward and reverse processes [20], providing a customized and adaptive approach to teleoperation tasks.
We conducted experiments in six different simulation environments using two types of end-effectors: a dexterous hand and a gripper. Additionally, we performed experiments on three real-world tasks to validate our findings. Evaluation results indicate that our proposed system significantly enhances data collection efficiency, increasing the collection success rate by 30% and nearly doubling the collection speed. Additionally, data collected in shared autonomy mode is as effective for downstream tasks and models as data collected directly from the teleoperation system, demonstrating comparable validity. Our main contributions are summarized as follows:
  • We study how to reduce human adaptation in teleoperation data collection and propose a human-agent joint learning paradigm.
  • We build a system that fosters concurrent development between the human operator and assistive agent, which not only streamlines the learning process but also expedites the robot’s ability to perform robot manipulations autonomously.
  • Conducting experiments to demonstrate the efficiency and effectiveness of our proposed system. Our system achieved significant performance improvements, including a $\mathit{30}\%$ increase in data collection success rate and nearly double the collection speed. We also deployed our system in a real-world environment and achieved significant results.
2 Related Works
Teleoperation System. Data has always been a crucial foundation, and robots are no exception. Teleoperation serves as a significant source for collecting robot data [7, 21, 22, 23, 24, 25]. Some works achieve teleoperation through wearable devices [1, 2, 3, 4, 26], and vision-based teleoperation systems offer a low-cost and easily developed alternative [5, 6, 27, 28]. For instance, Li et al. [28] utilizes neural networks for markerless vision-based teleoperation of dexterous robotic hands from depth images. Handa et al. [5] set up a vision-based teleoperation system to control the Allegro Hand, accomplishing various contact-rich manipulation tasks in the real world. Recently, Qin et al. [6] introduced AnyTeleop, a unified teleoperation system designed to accommodate various arms, hands, realities, and camera setups within a singular framework. In this paper, we introduce a joint learning paradigm to assist teleoperation by sharing control between the human operator and a learning-based agent, aiming to improve the efficiency of the teleoperation process.
Human Robot Cooperation. Collecting fine-grained human demonstration data for robotic manipulation is an effective but labor-intensive and time-consuming way to enable robots to complete a wide range of tasks [29, 30]. Previous work uses shared autonomy to assist people with disability in performing tasks by arbitrating human inputs and robot actions [31]. Many of the shared autonomy algorithms aim to estimate human intents from a set of pre-defined goals [32, 33, 34, 35] or by mapping low-dimension control input to high-dimension robot actions [31, 36]. In this work, we introduce a system that enables shared control between the human and assistive agent to facilitate the process of data collection and robot learning.
3 Technical Approach
The proposed system enables human operators to control the robot using a teleoperation system to gather training data (Sec. 3.1). Subsequently, utilizing the collected data, we train an agent (Sec. 3.2) to establish shared control between the human operator and the learned agent, thereby enhancing the efficiency of the data collection process (Sec. 3.3). Similar to the concept of “bootstrapping”, as more data accumulates, our system raises the control ratio of the learned agent, thereby reducing the effort required from human operators. This, in turn, enables us to collect even more data and continue improving the system iteratively. Moreover, we offer the option to transition the shared control agent to full autonomy once sufficient data is acquired (Sec. 3.4).
3.1 Teleoperation System.
Our pipeline initially captures the raw sensory signal $\mathcal{I}$. Human hand pose $\mathcal{P}^{h}\in\mathbb{R}^{20\times 3}$ can be obtained from the captured signal using off-the-shelf 3D hand pose estimation [10, 11, 13]. The pose $\mathcal{P}^{h}$ consists of the positions of $20$ keypoints. Then, employing an inverse kinematic function $f_{\textit{IK}}$, we compute the action of the robot $a\in\mathbb{R}^{m}$: $a=f_{\textit{IK}}(\mathcal{P}^{h}_{t},\mathcal{P}^{h}_{t+1})$, where it is calculated upon the change in the hand pose. Given this teleoperation system, the human operator will move the hand to produce a sequence of hand poses $\{\mathcal{P}^{h}_{i}\}_{i=0}^{T}$ to teleoperate the robot with an action sequence $\{a_{i}\}_{i=0}^{T}$ to achieve the task $\mathcal{T}$. The trajectory $\{(s_{i},a_{i})\}_{i=0}^{T}$ is the collected human demonstration data, where $s\in\mathbb{R}^{n}$ (here $n=18$) is the robot state, could be used for downstream tasks.
3.2 Diffusion-Model-Based Assistive Agent.
After getting the data, we train a diffusion-model-based assistive agent to learn how to assist the human in collecting data in a shared control manner. We follow the Denoising Diffusion Probabilistic Model (DDPM) [18] training paradigm to construct the diffusion-model-based assist agent. The forward process of the Diffusion Model can be regarded as adding Gaussian noise to the data $x^{0}$ according to a variance schedule $\beta_{1:K}$ by
$x_{k}=\sqrt{\alpha_{k}}x_{k-1}+\sqrt{1-\alpha_{k}}\epsilon,$ (1)
where $\epsilon\sim\mathcal{N}(\mathbf{0,I}),\alpha_{k}=1-\beta_{k}$. DDPM models the output generation as a denoising process (Stochastic Langevin Dynamics). A line of works [19, 17, 8, 37] use diffusion model to generate the action for agents: given $x^{K}$ sampled from Gaussian noise $\mathcal{N}(\mathbf{0,I})$, it utilizes a parameterized diffusion process to model how $x^{K}$ is denoised in order to get noise-free action $x^{0}$ by
$p_{\theta}(x^{0})=\int p(x^{K})\prod_{k=1}^{K}p_{\theta}(x^{k-1}|x^{k})\mathrm{d}x^{1:K},$ (2)
where $p_{\theta}(x^{k-1}|x^{k})=\mathcal{N}(\mu_{\theta}(x^{k},k),\Sigma(x^{k},k))$ is usually referred as reverse process. Luo [38] shows that $p_{\theta}(x^{t-1}|x^{k})$ becomes tractable when conditioned on $x_{0}$ and Eq. 2 can be reformulated as minimizing the error in the noise prediction. Ho et al. [18] simplify the training loss function as
$\mathcal{L}:=\mathbb{E}_{k,\boldsymbol{x}_{0},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{I})}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{k}(\boldsymbol{x}_{0},\boldsymbol{\epsilon}),k)\|_{2}^{2}\right],$ (3)
where step $k$ is sampled uniformly as $k\in[1,K]$, $\boldsymbol{\epsilon}_{\theta}$ is the noise prediction model. During the inference phase, we can generate $x_{0}$ by recursively sample $\boldsymbol{z}\sim\mathcal{N}(\mathbf{0},\boldsymbol{I})$:
$x_{k-1}=\mu_{\theta}(x_{k},k)+\sigma_{k}\boldsymbol{z}.$ (4)
Similar to [8, 20], with the collected trajectory $\{(s_{i},a_{i})\}_{i=0}^{T}$, we aim to train an agent to imitate the trajectory, accomplishing a specific task $\mathcal{T}$. Therefore, we utilize DDPM to capture the conditional distribution of $p(a|s)$ and the training loss in Eq. 3 shall be modified as
$\mathcal{L}:=\mathbb{E}_{k,(s_{i},a_{i}),\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{I})}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}({a}_{i}+\boldsymbol{\epsilon},s_{i},k)\|_{2}^{2}\right].$ (5)
At an abstract level, the diffusion-model-based assist agent, noted as $f(\cdot|\cdot)$, is provided with the state $s$, denoising step number $k$, and a noise action $a^{k}$, which could be an imperfect action gathered from the teleoperation system or sampled from a Gaussian distribution, to predict the desired action
$a=f(a^{k}|s,k).$ (6)
Note that during the experiment, we found that adding slight noise to $s_{i}$ during training will give a better result. We assume doing so will augment the dataset.
3.3 Data Collection with Shared Control
Figure 2: To achieve shared control between the human and agent, we blend the action from the human operator $a^{h}$ using the forward and reverse process. The parameter $\gamma$ governs the control ratio, where a lower $\gamma$ results in the action better aligning with the human operator’s intention. In contrast, a higher $\gamma$ allows the learned agent to exert more influence over the blended action.
During data collection, the proposed system offers the option to control the robot in a shared control mode rather than directly applying the collected action $a^{h}$ from the teleoperation system. The classical shared autonomy method is achieved through the equation [32]:
$a^{s}=\gamma a^{h}+(1-\gamma)a^{r},$ (7)
where $a^{r}$ is generated by the learned agent. Considering that the agent operates as a diffusion policy (Fig. 2), we blend the action from the human with the forward and reverse processes. Given action $a^{h}$, a forward process diffuses the action as follows: $a^{k}=a^{h}+\epsilon^{k}.$ Subsequently, a reverse process denoises the action $a^{k}$:
$a^{s}=f(a^{k}|s,k).$ (8)
By applying action $a$, the control of the robot is shared between the human and the diffusion-model-based assistive agent. We can adjust the control ratio $\gamma=k/K$ between the human operator and the diffusion-model-based assistive agent by varying $k$. When $\gamma=0$, the action $a^{s}$ represents the teleoperation action $a^{h}$, which is the dexterous robot directly controlled by a human operator. As $\gamma$ approaches $1.0$, the action $a^{s}$ transitions to full autonomy $a^{r}$. A higher $\gamma$ value indicates a higher level of autonomy, allowing the learning-based agent more control rights to stabilize and direct the dexterous hand. This leads to a reduced human workload during the data collection process.
3.4 Intergrating Data Collection and Manipulation Learning.
We outline the overall process in Algo. 1. The assistive agent is trained in four steps as follows:
Step 1. Initially, we collect a dataset for pre-training agent $f$ under full manual control by human operators, i.e., with the control ratio $\gamma=0$.
Step 2. Given the initial dataset, we train a less capable assistive agent to aid in further data collection. The training process has been formulated in Eq. 5 and Eq. 6, where a neural network $\epsilon_{\theta}$ is trained to predict noise $\epsilon$ out of the noisy action $a^{k}$.
Step 3. The trained agent assists in a second data collection round, aiming for higher efficiency and success. We refine the agent using data from both rounds to enhance its performance. This cycle repeats until the agent achieves full autonomy and the required data volume is collected.
Algorithm 1 Overall Process
1:The human operator $\mathcal{H}$;
2:The collected dataset $\mathcal{D}$; the learned agent $f$; control ratio $\gamma$;
3:Initialization: $\mathcal{D}\leftarrow\emptyset,\gamma\leftarrow 0$;
4:while $|\mathcal{D}|$ is small do $\triangleright$ not enough data is collected
5:     $\mathcal{H}$ collects data $d$ under $f$’s help; $\triangleright$ with proper control ratio $\gamma$, details shown in Fig. 2
6:     if $d$ is valid then
7:         $\mathcal{D}\leftarrow\{d\}\cup\mathcal{D}$;
8:     Finetune $f$ with $\mathcal{D}$;
9:     Raise the agent’s control ratio $\gamma$;
10:return $\mathcal{D}$ and $f$;
Currently, the adjustment of the control ratio $\gamma$ is guided by intuitive assessments and careful monitoring of success rates. When an agent, trained with recently gathered data, exhibits a marked improvement in success rate relative to its prior performance, this enhancement is taken as a cue to fine-tune $\gamma$, thus refining control effectiveness. The assessment of the agent’s performance, followed by adjustments to $\gamma$, relies on empirical data and the agent’s performance in practice. This method ensures that decisions regarding the adjustment of $\gamma$ are firmly rooted in data, thus harmonizing the control strategy with the observed outcomes. It is posited that an agent reaches full autonomy when it achieves a success rate that aligns with or surpasses current state-of-the-art benchmarks.
4 Experiments
Tasks. We adopt six multi-stage manipulation tasks (Fig. 8). Pick-and-Place aims at picking an object on the table and placing it into a container. Articulated-Manipulation’s objective for the dexterous hand is to grasp and unscrew a door handle to open it, while for the gripper, it is to grab a drawer handle and pull the drawer open. Push-cube requires the robot to push the cube to the target position. Tool-Use aims at picking a hammer and using it to drive a nail into a board.
Efficiency of Data Collection. Our proposed system leverages shared control between human operators and learned agents to enhance the efficiency of data collection. To learn how the assistant agent could improve the data collection process, we conducted a user study.
Figure 8: Overview of six task settings and their task flow for Pick-and-Place (left), Articulated-Manipulation (middle), Gripper-Push (upper-right) and Dexterous-Tool-Use (bottom-right).
In the user study, 10 human operators participate, collecting data under two modes: one where control is shared between the operator and the learned agent (w/ Ours), and the other where control is directly by the operator alone (w/o Ours). Each participant is instructed to collect as much data as possible within 3 minutes under two different modes for 3 dexterous hand tasks. Three metrics are evaluated: Success Rate (Percent) indicates the percentage of attempts where data collection was successful. Horizon Length (Steps per Sample) measures the length of each collected trajectory, with a lower horizon length indicating smoother data collection. Collection Speed (Samples per Hour) refers to the number of successful trajectories that can be collected in one hour.
Pick-and-Place Door-Open Tool-Use
Success Horizon Collection Success Horizon Collection Success Horizon Collection
Rate $\uparrow$ Length $\downarrow$ Speed $\uparrow$ Rate $\uparrow$ Length $\downarrow$ Speed $\uparrow$ Rate $\uparrow$ Length $\downarrow$ Speed $\uparrow$
Group1 w/ Ours 86.96 219.01 320 87.11 142.29 460 66.50 232.17 200
w/o Ours 51.53 378.49 176 62.49 258.27 252 42.38 487.95 129
Group2 w/ Ours 94.06 214.16 324 80.29 134.16 424 55.55 275.71 172
w/o Ours 45.42 471.48 120 53.45 317.21 176 34.47 511.03 124
Table 1: User studies on three dexterous hand tasks.
Figure 9: Shared control process overview. The white one is the hand controlled purely by the human operator, while the cyan one is under shared control between the human and the assistive agent.
In Tab. 1, by sharing control between humans and learned agents, our system shows improvements in both success rate and collection speed, while the average horizon length of the collected trajectories is reduced. This suggests that our system enhances the efficiency of data collection by facilitating a process that is easier to succeed, faster, and more fluid in terms of trajectory smoothness. To ensure the fairness of the experiment wasn’t compromised, we equally divided the user group into two parts, Group 1 first collected data directly by themselves (w/o Ours) and then collected data with an assistive agent (w/ Ours), while the Group 2 reversed the order, first (w/ Ours) mode and then (w/o Ours) mode.
To gain deeper insight into how the learned agent assists the human operator, we visualize several keyframes from the data collection process of three dexterous hand tasks under shared control mode with $\gamma=0.4$. From Fig. 9, it is evident that human operators are not required to provide too precise control with the assistive agent facilitating shared control over the dexterous hand. Instead, they only need to convey high-level intentions, such as the direction of hand movement or finger grasp motions. In multi-stage tasks, like picking up a hammer and then using it to drive a nail, operators only need to provide a trigger action to guide the agent to transition from one sub-stage to the next. As a result, less effort and attention are required, making the data collection easier to execute successfully and speeding it up.
Quantitative Evaluation. When the learned agent shares control with users, the system effectively corrects imperfect human control signals to achieve specific tasks. Given the challenge of directly measuring the level of imperfection in user signals and the correction ability of our system, we simulate the human using a baseline agent trained with Behavior Cloning (BC). Additionally, we import noise to the agent’s control signal through the diffusion policy’s forward diffuse process.
In Fig. LABEL:fig:simulateduser, the x-axis represents the forward ratio, where a higher forward ratio corresponds to the lower quality of action provided by the simulated operator and a higher control ratio of the learned agent on the dexterous hand. The graph illustrates that with limited data availability, the agent can assist the simulated operator more effectively. As the agent accesses and trains with more data, its ability to correct actions improves. The results show that our system gradually diminishes the demand for the operator’s attention and effort, thereby enhancing the efficiency of data collection.
Furthermore, once sufficient data is collected and the assistive agent is trained, it can transition into full autonomy mode by setting $\gamma$ to 1.0 and denoising random actions from the Gaussian distribution noise. Across three different dexterous manipulation tasks, we can achieve success rates of 0.76, 0.78, and 0.89, indicating that the assistive agent can effectively transform into an automated dexterous manipulation agent.
From our experiments, we have observed that the assistive agent significantly aids human operators in managing fine control at the low level, especially in scenarios where accurate observation by humans is challenging, complicating effective action control. For instance, tasks such as grasping an egg or moving a hammer present visual challenges. It can be difficult to visually confirm whether the egg is securely grasped or if there’s a risk of it being dropped. This uncertainty makes it hard for human operators to react promptly to sudden changes. However, within our proposed joint learning framework, human operators are primarily required to focus on high-level intentions and task planning during data collection, while the assistive agent manages the detailed low-level actions. This division of labor significantly reduces the burden on human operators by clearly separating strategic planning from execution tasks, streamlining the collaboration between humans and machines.
Dexterous Pick-and-Place Articulated-Manipulation Tool-Use
[…]
BC 0.30 0.50 0.22 0.57 0.39 0.40
BC-RNN 0.54 0.67 0.47 0.50 0.27 0.25
DP 0.73 0.76 0.77 0.78 0.88 0.89
Parallel Pick-and-Place Articulated-Manipulation Push-cube
[…]
BC 0.42 0.44 0.35 0.37 0.88 0.85
BC-RNN 0.39 0.36 0.71 0.73 0.59 0.67
DP 0.51 0.60 0.42 0.67 0.83 0.82
Table 2: Data quality on downstream tasks.
Dexterous Tool-Use Gripper Push-cube
BC DP BC DP
$\mathcal{H}$ 0.29 0.45 0.23 0.42
$\mathcal{H}$ + $\mathcal{H}$ 0.28 0.67 0.37 0.78
$\mathcal{H}$ + $\mathcal{H}$ 0.28 0.82 0.51 0.67
$\mathcal{H}$ + $\mathcal{H}$ 0.39 0.88 0.88 0.83
[…]
Table 3: Tool-Use and Push-cube task success rate under increasing data.
Data Quality on Downstream Task. In this section, we illustrate that collecting data under shared control does not compromise the quality of the data. We gather dexterous hand and gripper manipulation demonstrations via the proposed system in two modes: fully controlling the robots by a human ($\mathcal{H}$) and sharing control ($\mathcal{S}$) between the human operator and the learned assistive agent. And utilize these data to train different kinds of agents, like BC, BC-RNN [7], and Diffusion Policy (DP) [8].
In Tab. 2, compared to directly collecting human demonstrations from the expert human operator, who can achieve success rates and efficiency comparable to those with agent assistance, the data collected by sharing control between the human and the assistive agent can achieve comparable or even surprisingly better results with BC and BC-RNN. Their results are comparable with DP, possibly as DP can better fit the tasks, which is in line with [8].
In Tab. 3, we compare the effects of using different sets of data to train BC and DP. We can find that utilizing more data collected under the shared control mode leads to comparable performance on the tool-use task. This verifies that the new data contributes significantly to policy learning and can achieve a similar effect compared to the data from human experts but at a much lower cost. These results indicate that the data collected under the proposed paradigm have sufficient quality and efficiency for downstream tasks.
Figure 12: Real world setting. 1. Pick-and-Place: (a) Hardware setup. The robot gripper should pick the red pot up (b-c) and place it onto the black induction cooker (d). 2. Articulated-Manipulation: use a gripper to open the drawer. 3. Push-cube: use a gripper to push the cube across the line.
Success Rate $\uparrow$ Horizon Length $\downarrow$ Collection Speed $\uparrow$
w/ Ours 0.79 18.72 151
w/o Ours 0.70 21.54 121
Table 4: Real world parallel gripper Pick-and-Place task user study.
Pick-and-Place Articulated-Manipulation Push-cube
[…]
BC 13 / 20 14 / 20 18 / 20 19 / 20 15 / 20 15 / 20
DP 11 / 20 12 / 20 16 / 20 12 / 20 15 / 20 13 / 20
Table 5: Real world parallel gripper experiments of data quality.
Real World Experiment. To better evaluate our system, we further conduct real-world experiments. Three tasks are adopted: Pick-and-Place, Articulated-Manipulation, and Push-cube in Fig. 15. Following the same rules as Sec. 4, four human volunteers are invited to participate in the user study to collect data under two modes: one where control is shared between the human operator and the learned agent (w/ Ours), and the other where control is directly by the human operator alone (w/o Ours). Our proposed system achieves significant improvements in success rate and collection speed by sharing control between human operators and learned agents, as demonstrated in Tab. 4. Additionally, data gathered under our proposed joint learning shared control mode yield performance on the three tasks that are comparable to those pure human datasets using BC and DP, further substantiated by the results presented in Tab. 5.
5 Conclusion
In this paper, we introduce a novel human-agent joint learning paradigm that enables simultaneous human demonstration collection and robot manipulation teaching. This approach allows the human operator to share control with a diffusion-model-based assistive agent within a vision-based teleoperation system to control multiple robot end-effectors such as grippers and dexterous hands. Given our paradigm, the human operator can reduce the effort spent on data collection and adjust the control ratio between the human and agent based on different scenarios. Our system offers a more efficient and flexible solution for data collection and robot manipulation learning via teleoperation.
Acknowledgments
This work is supported in part by the National Natural Science Foundation of China under Grants 62306175.

### v3

@@
 Employing a teleoperation system for gathering demonstrations offers the potential for more efficient learning of robot manipulation.
-However, teleoperating a robot arm equipped with a dexterous hand or gripper, via a teleoperation system poses significant challenges due to its high dimensionality, complex motions, and differences in physiological structure.
-In this study, we introduce a novel system for joint learning between human operators and robots, that enables human operators to share control of a robot end-effector with a learned assistive agent, facilitating simultaneous human demonstration collection and robot manipulation teaching.
-In this setup, as data accumulates, the assistive agent gradually learns.
+However, teleoperating a robot arm equipped with a dexterous hand or gripper, via a teleoperation system presents inherent challenges due to the task’s high dimensionality, complexity of motion, and differences between physiological structures.
+In this study, we introduce a novel system for joint learning between human operators and robots, that enables human operators to share control of a robot end-effector with a learned assistive agent, simplifies the data collection process, and facilitates simultaneous human demonstration collection and robot manipulation training.
+As data accumulates, the assistive agent gradually learns.
 Consequently, less human effort and attention are required, enhancing the efficiency of the data collection process.
@@
 Through user studies and quantitative evaluations, it is evident that the proposed system could enhance data collection efficiency and reduce the need for human adaptation while ensuring the collected data is of sufficient quality for downstream tasks.
-Videos are available at https://norweig1an.github.io/human-agent-joint-learning.github.io/.
-Keywords: Robotic Teleoperation, Robot Manipulation, Imitation Learning
-footnotetext: * denotes equal contribution, †denotes corresponding author.
-1 Introduction
-The long-term vision in the field of robot learning has been to enable robots to perform diverse tasks at a human level in the physical world.
-Recently, significant progress has been made toward this goal by learning robot manipulation policies from demonstrations.
-Previous studies have utilized teleoperation systems [1, 2, 3, 4, 5, 6] to collect human demonstrations, and learning-based policies [7, 8, 9] have been formulated using the gathered data.
-Despite the notable advancements, several challenges still need to be addressed.
-For example, in vision-based teleoperation systems, even with state-of-the-art 3D hand pose estimation algorithms [10, 11, 12, 13], errors persist that significantly affect the teleoperation.
-Additionally, discrepancies between the structures of human hands and robot end-effectors, along with the lack of haptic feedback during contact-rich manipulation, also pose challenges.
-As a result, current teleoperation systems require human operators to practice extensively to adapt to these differences and gather the necessary data.
-This means that many human adaptations are essential.
-Furthermore, to meet the data requirements of robotic systems, humans need to collect large amounts of data, making this process very burdensome.
-Figure 1: Traditional frameworks typically separate human and agent training, requiring operators first to learn the task environment before data collection.
+For more details, please refer to our webpage https://norweig1an.github.io/HAJL.github.io/.
+{strip}[Uncaptioned image]
+Figure 1: Human-agent joint learning overview.
+Traditional frameworks typically separate human and agent training, requiring operators first to learn the task environment before data collection.
 This often leads to inefficiencies due to delayed and insufficient data gathering.
@@
 This enables simultaneous development and adapts the agents to human operation more effectively, enhancing overall efficiency and promoting better collaboration between humans and machines allowing for human effortless adaptation data collection.
-Naturally, a question was raised: in data-collection, how to make human adaptation less or even free while keeping the data quality?
+I INTRODUCTION
+In recent years, significant progress has been made on learning robot manipulation policies from demonstrations.
+Previous studies have utilized teleoperation systems [1, 2, 3, 4, 5, 6] to collect human demonstrations, and learning-based policies [7, 8, 9] have been formulated using the gathered data.
+Despite the notable advancements, several challenges still need to be addressed.
+For example, in vision-based teleoperation systems, even with state-of-the-art 3D hand pose estimation algorithms [10, 11, 12, 13], errors persist that significantly degrade the teleoperation.
+Additionally, discrepancies between the structures of human hands and robot end-effectors, along with the lack of haptic feedback during contact-rich manipulation, also pose challenges.
+As a result, current teleoperation systems demand substantial human effort, in addition, collecting high-quality datasets remains a challenging and labor-intensive task in many scenarios.
+Naturally, a question was raised: in data-collection, how to make human effort less while improving the data quality?
 Here, we aim to address this question and argue that human-agent joint learning can help.
 That said, an effective and efficient teleoperation system should be designed to preferentially capture the operator’s intentions for directing a robot end effector and pose the main frame, while concurrently enabling an autonomous agent to help us ensure motion stability and interpolate the details.
-To this end, we propose a framework that achieves shared control between the human and a learned assistive agent.
+To this end, we propose a framework that achieves shared control between the human and a learned assistive agent during data collection.
 As shown in Fig.
-1, our human effortless adaptation framework seeks to balance human needs and system performance, aiming to directly enhance the efficiency and quality of data collection by reducing the time and effort in human adaptation.
-Given our human-agent joint learning approach, we allow the data acquisition agent to grow and learn along with the human operator, reducing the learning burden of the human operator throughout the data collection task.
-Inspired by shared autonomy [14, 15, 16], we introduce a novel teleoperation system that enables collaboration between humans and learning-based agents to control a robot jointly.
+1, our human-agent joint learning framework seeks to integrate the data collection and policy learning to enhance the efficiency of the whole process, reducing human effort, and improving the data quality.
+Given our human-agent joint learning approach, we allow the data acquisition agent to grow and learn along with the human operator.
+Inspired by shared autonomy [14, 15, 16], we introduce a novel teleoperation system that enables collaboration between humans and learning-based agents to control a robot jointly during the data collection and learning process.
 In particular, our proposed system provides the flexibility to adjust a “control ratio” between the human operator and a learning-based agent.
-A lower control ratio, in the beginning, signifies a greater emphasis on humans to teach the agent the finer-grained knowledge under the structure of human intention and principal actions.
-As the agent’s learning improves, a higher ratio later indicates increased autonomy from the learned agent to replace the human effort to “inpaint” the whole process given only human intention and principal actions.
-Additionally, once a sufficient amount of data is gathered, we have the option to transition the shared autonomy agent to full autonomy by reducing the human control ratio to zero.
-We implement our framework as a joint learning system, which facilitates a human operator’s intention and an assistive agent’s execution of teleoperation tasks.
-For human operators, their inputs consist of intuitive control actions based on visual feedback and past cognitive experience.
-For the agent, their inputs are derived from a combination of sensors, data streams, and possibly pre-processed information.
-This can include visual data from cameras [1], tactile data from force sensors [3], and any additional context provided by the system [6], such as object recognition or environmental mapping.
-The agent processes these inputs using algorithms designed to interpret the task requirements and generate actions to assist the human operator, utilizing a diffusion model [8, 17, 18, 19] as the backbone enables us to adjust the control ratio by modifying the step number of the forward and reverse processes [20], providing a customized and adaptive approach to teleoperation tasks.
+A lower control ratio, in the beginning, emphasizes the human’s role in teaching the agent finer-grained knowledge under the structure of human intention and principal actions.
+As the agent’s learning improves, a higher ratio indicates greater autonomy from the learned agent to replace the human effort to “inpaint” the whole process given only human intention and principal actions.
+With the proposed system, the human effort will be reduced due to the shared control during data collection.
+Additionally, the agent learning process is integrated with the data collection, improving the efficiency of the whole process.
+In addition, the quality of the collected data is also improved, benefiting different kinds of downstream tasks.
 We conducted experiments in six different simulation environments using two types of end-effectors: a dexterous hand and a gripper.
@@
 Our main contributions are summarized as follows:
-  • We study how to reduce human adaptation in teleoperation data collection and propose a human-agent joint learning paradigm.
+  • We study how to reduce human adaptation while keeping data quality in teleoperation data collection and propose a human-agent joint learning paradigm.
   • We build a system that fosters concurrent development between the human operator and assistive agent, which not only streamlines the learning process but also expedites the robot’s ability to perform robot manipulations autonomously.
-  • Conducting experiments to demonstrate the efficiency and effectiveness of our proposed system.
-Our system achieved significant performance improvements, including a $\mathit{30}\%$ increase in data collection success rate and nearly double the collection speed.
-We also deployed our system in a real-world environment and achieved significant results.
-2 Related Works
-Teleoperation System.
+  • Conducting both simulation and real-world experiments to demonstrate the efficiency and effectiveness of our proposed system.
+Our system achieved significant performance improvements, including a 30% increase in data collection success rate and double the collection speed.
+II RELATED WORKS
+Teleoperation for Data Collection.
 Data has always been a crucial foundation, and robots are no exception.
-Teleoperation serves as a significant source for collecting robot data [7, 21, 22, 23, 24, 25].
+Teleoperation serves as a significant source for collecting robot data [7, 17, 18, 19, 20, 21, 22].
 Some works achieve teleoperation through wearable devices [1, 2, 3, 4, 23], and vision-based teleoperation systems offer a low-cost and easily developed alternative [5, 6, 24, 25].
-For instance, Li et al.
-[28] utilizes neural networks for markerless vision-based teleoperation of dexterous robotic hands from depth images.
-Handa et al.
+For instance, [25] utilizes neural networks for markerless vision-based teleoperation of dexterous robotic hands from depth images.
 [5] set up a vision-based teleoperation system to control the Allegro Hand, accomplishing various contact-rich manipulation tasks in the real world.
-Recently, Qin et al.
-[6] introduced AnyTeleop, a unified teleoperation system designed to accommodate various arms, hands, realities, and camera setups within a singular framework.
-In this paper, we introduce a joint learning paradigm to assist teleoperation by sharing control between the human operator and a learning-based agent, aiming to improve the efficiency of the teleoperation process.
-Human Robot Cooperation.
+Recently, [6] introduced AnyTeleop, a unified teleoperation system designed to accommodate various arms, hands, realities, and camera setups within a singular framework.
+In this paper, we introduce a joint learning paradigm to assist teleoperation by sharing control between the human operator and a learning-based agent, improving the efficiency of data collection using teleoperation.
+Interactive robot learning.
 Collecting fine-grained human demonstration data for robotic manipulation is an effective but labor-intensive and time-consuming way to enable robots to complete a wide range of tasks [26, 27].
 Previous work uses shared autonomy to assist people with disability in performing tasks by arbitrating human inputs and robot actions [28].
-Many of the shared autonomy algorithms aim to estimate human intents from a set of pre-defined goals [32, 33, 34, 35] or by mapping low-dimension control input to high-dimension robot actions [31, 36].
-In this work, we introduce a system that enables shared control between the human and assistive agent to facilitate the process of data collection and robot learning.
-3 Technical Approach
-The proposed system enables human operators to control the robot using a teleoperation system to gather training data (Sec.
-3.1).
-Subsequently, utilizing the collected data, we train an agent (Sec.
-3.2) to establish shared control between the human operator and the learned agent, thereby enhancing the efficiency of the data collection process (Sec.
-3.3).
-Similar to the concept of “bootstrapping”, as more data accumulates, our system raises the control ratio of the learned agent, thereby reducing the effort required from human operators.
-This, in turn, enables us to collect even more data and continue improving the system iteratively.
-Moreover, we offer the option to transition the shared control agent to full autonomy once sufficient data is acquired (Sec.
-3.4).
-3.1 Teleoperation System.
-Our pipeline initially captures the raw sensory signal $\mathcal{I}$.
-Human hand pose $\mathcal{P}^{h}\in\mathbb{R}^{20\times 3}$ can be obtained from the captured signal using off-the-shelf 3D hand pose estimation [10, 11, 13].
-The pose $\mathcal{P}^{h}$ consists of the positions of $20$ keypoints.
-Then, employing an inverse kinematic function $f_{\textit{IK}}$, we compute the action of the robot $a\in\mathbb{R}^{m}$: $a=f_{\textit{IK}}(\mathcal{P}^{h}_{t},\mathcal{P}^{h}_{t+1})$, where it is calculated upon the change in the hand pose.
-Given this teleoperation system, the human operator will move the hand to produce a sequence of hand poses $\{\mathcal{P}^{h}_{i}\}_{i=0}^{T}$ to teleoperate the robot with an action sequence $\{a_{i}\}_{i=0}^{T}$ to achieve the task $\mathcal{T}$.
-The trajectory $\{(s_{i},a_{i})\}_{i=0}^{T}$ is the collected human demonstration data, where $s\in\mathbb{R}^{n}$ (here $n=18$) is the robot state, could be used for downstream tasks.
-3.2 Diffusion-Model-Based Assistive Agent.
-After getting the data, we train a diffusion-model-based assistive agent to learn how to assist the human in collecting data in a shared control manner.
-We follow the Denoising Diffusion Probabilistic Model (DDPM) [18] training paradigm to construct the diffusion-model-based assist agent.
+Many of the shared autonomy algorithms aim to estimate human intents from a set of pre-defined goals [29, 30, 31, 32], using clothoid curves to parametrize the state and control [33] or by mapping low-dimension control input to high-dimension robot actions [28, 34].
+In this work, we introduce a system that integrates the agent’s learning process with data collection, facilitating both data collection and robot learning.
+III PROPOSED METHOD
+The primary contribution of this work is the development of a novel and highly efficient data collection method.
+To achieve this, the system is designed in two key stages: first, the proposed system allows human operators to control the robot via a teleoperation system to gather an initial but insufficient training dataset, which serves as the foundation for the second stage (Sec.
+III-B).
+Second, using these data, we train a diffusion-model-based assistive agent (Sec.
+III-C) to establish shared control between the human operator and the agent, thereby improving the efficiency of the data collection process (Sec.
+III-D).
+This approach mirrors the concept of “bootstrapping” [35], where, as more data is accumulated, the system progressively reduces the effort required from human operators, facilitating further data collection and iterative system improvement.
+Additionally, once sufficient data has been gathered, the system offers the option to transition the shared control agent to full autonomy.
+III-A Preliminary
+To get a learned agent in Sec.
+III-C, enabling human-agent joint learning, we follow the Denoising Diffusion Probabilistic Model (DDPM) [36] training paradigm.
+Here we first briefly introduce the DDPM algorithm.
 The forward process of the Diffusion Model can be regarded as adding Gaussian noise to the data $x^{0}$ according to a variance schedule $\beta_{1:K}$ by
@@
 where $p_{\theta}(x^{k-1}|x^{k})=\mathcal{N}(\mu_{\theta}(x^{k},k),\Sigma(x^{k},k))$ is usually referred as reverse process.
-Luo [38] shows that $p_{\theta}(x^{t-1}|x^{k})$ becomes tractable when conditioned on $x_{0}$ and Eq.
+[40] shows that $p_{\theta}(x^{t-1}|x^{k})$ becomes tractable when conditioned on $x_{0}$ and Eq.
 2 can be reformulated as minimizing the error in the noise prediction.
-Ho et al.
 [36] simplify the training loss function as
@@
 $\mathcal{L}:=\mathbb{E}_{k,(s_{i},a_{i}),\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{I})}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}({a}_{i}+\boldsymbol{\epsilon},s_{i},k)\|_{2}^{2}\right].$ (5)
+Figure 7: Simulation tasks overview.
+Here are six task settings and their task flow for Pick-and-Place (left), Articulated-Manipulation (middle), Gripper-Push (upper-right) and Dexterous-Tool-Use (bottom-right).
+III-B Teleoperation System.
+Our system initially captures the raw sensory signal $\mathcal{I}$.
+Human hand pose $\mathcal{P}^{h}$ can be obtained from the captured signal using off-the-shelf 3D hand pose estimation [10, 11, 13].
+The pose $\mathcal{P}^{h}$ consists of the positions of the human hand’s keypoints.
+Then, employing an inverse kinematic function $f_{\textit{IK}}$, we compute the action of the robot $a\in\mathbb{R}^{m}$, such that $a=f_{\textit{IK}}(\mathcal{P}^{h}_{t},\mathcal{P}^{h}_{t+1})$, where it is calculated upon the change in the hand pose.
+Given this teleoperation system, the human operator will move the hand to produce a sequence of hand poses $\{\mathcal{P}^{h}_{i}\}_{i=0}^{T}$ to teleoperate the robot with an action sequence $\{a_{i}\}_{i=0}^{T}$ to achieve the task $\mathcal{T}$.
+The human collected demonstration trajectory $\{(s_{i},a_{i})\}_{i=0}^{T}$, where $s\in\mathbb{R}^{n}$ is the robot state, could be used for downstream tasks.
+III-C Diffusion-Model-Based Assistive Agent.
+After collecting data via teleoperation mentioned in Sec. III-B, we train a diffusion-model-based assistive agent to learn how to assist humans in collecting data in a shared control manner.
 At an abstract level, the diffusion-model-based assist agent, noted as $f(\cdot|\cdot)$, is provided with the state $s$, denoising step number $k$, and a noise action $a^{k}$, which could be an imperfect action gathered from the teleoperation system or sampled from a Gaussian distribution, to predict the desired action
 $a=f(a^{k}|s,k).$ (6)
-Note that during the experiment, we found that adding slight noise to $s_{i}$ during training will give a better result. We assume doing so will augment the dataset.
-3.3 Data Collection with Shared Control
-Figure 2: To achieve shared control between the human and agent, we blend the action from the human operator $a^{h}$ using the forward and reverse process.
+Figure 8: Diffusion based shared control.
+To achieve shared control between the human and agent, we blend the action from the human operator $a^{h}$ using the forward and reverse process.
 The parameter $\gamma$ governs the control ratio, where a lower $\gamma$ results in the action better aligning with the human operator’s intention.
@@
 During data collection, the proposed system offers the option to control the robot in a shared control mode rather than directly applying the collected action $a^{h}$ from the teleoperation system.
+This leads to a reduced human workload during the data collection process.
 The classical shared autonomy method is achieved through the equation [29]:
@@
 where $a^{r}$ is generated by the learned agent.
-Considering that the agent operates as a diffusion policy (Fig.
-2), we blend the action from the human with the forward and reverse processes.
+However, considering that the agent operates as a diffusion policy (Fig.
+8), we blend the action from the human with the forward and reverse processes.
 Given action $a^{h}$, a forward process diffuses the action as follows: $a^{k}=a^{h}+\epsilon^{k}.$ Subsequently, a reverse process denoises the action $a^{k}$:
@@
 A higher $\gamma$ value indicates a higher level of autonomy, allowing the learning-based agent more control rights to stabilize and direct the dexterous hand.
-This leads to a reduced human workload during the data collection process.
-3.4 Intergrating Data Collection and Manipulation Learning.
-We outline the overall process in Algo. 1. The assistive agent is trained in four steps as follows:
+Algorithm 1 Overall Process
+1:The human operator $\mathcal{H}$;
+2:The collected dataset $\mathcal{D}$; assistive agent $f$; control ratio $\gamma$;
+3:Initialization: $\mathcal{D}\leftarrow\emptyset,\gamma\leftarrow 0$;
+4:while $|\mathcal{D}|$ is small do $\triangleright$ not enough data is collected
+5:     $\mathcal{H}$ collects data $d$ under $f$’s help; $\triangleright$ see III-D2 for control ratio adjustment
+6:     if $d$ is valid then
+7:         $\mathcal{D}\leftarrow\{d\}\cup\mathcal{D}$;
+8:     end if
+9:     Finetune $f$ with $\mathcal{D}$;
+10:end while
+11:return $\mathcal{D}$ and $f$;
+III-D Integrating Data Collection and Manipulation Learning.
+In this section, we show how to integrate data collection and manipulation learning into a unified framework that progressively reduces human effort and enhances robot autonomy.
+III-D1 Detailed Algorithm Explanation
+We outline the overall process in Algo. 1. The assistive agent is trained in three steps as follows:
 Step 1. Initially, we collect a dataset for pre-training agent $f$ under full manual control by human operators, i.e., with the control ratio $\gamma=0$.
 Step 2.
-Given the initial dataset, we train a less capable assistive agent to aid in further data collection.
+Given the initial dataset, we train a relatively low performance assistive agent to aid in further data collection.
 The training process has been formulated in Eq.
@@
 The trained agent assists in a second data collection round, aiming for higher efficiency and success.
-We refine the agent using data from both rounds to enhance its performance.
+We then refine the agent using data from both rounds to enhance its performance.
 This cycle repeats until the agent achieves full autonomy and the required data volume is collected.
-Algorithm 1 Overall Process
-1:The human operator $\mathcal{H}$;
-2:The collected dataset $\mathcal{D}$; the learned agent $f$; control ratio $\gamma$;
-3:Initialization: $\mathcal{D}\leftarrow\emptyset,\gamma\leftarrow 0$;
-4:while $|\mathcal{D}|$ is small do $\triangleright$ not enough data is collected
-5:     $\mathcal{H}$ collects data $d$ under $f$’s help; $\triangleright$ with proper control ratio $\gamma$, details shown in Fig. 2
-6:     if $d$ is valid then
-7:         $\mathcal{D}\leftarrow\{d\}\cup\mathcal{D}$;
-8:     Finetune $f$ with $\mathcal{D}$;
-9:     Raise the agent’s control ratio $\gamma$;
-10:return $\mathcal{D}$ and $f$;
-Currently, the adjustment of the control ratio $\gamma$ is guided by intuitive assessments and careful monitoring of success rates.
-When an agent, trained with recently gathered data, exhibits a marked improvement in success rate relative to its prior performance, this enhancement is taken as a cue to fine-tune $\gamma$, thus refining control effectiveness.
-The assessment of the agent’s performance, followed by adjustments to $\gamma$, relies on empirical data and the agent’s performance in practice.
-This method ensures that decisions regarding the adjustment of $\gamma$ are firmly rooted in data, thus harmonizing the control strategy with the observed outcomes.
-It is posited that an agent reaches full autonomy when it achieves a success rate that aligns with or surpasses current state-of-the-art benchmarks.
-4 Experiments
-Tasks.
+III-D2 Control Ratio Adjustment
+For each data collection, we offer users two options to adjust the control ratio $\gamma$: (1) Users can empirically adjust $\gamma$ based on their needs.
+(2) Alternatively, set $\gamma=\frac{1}{2}(1+\cos\theta)$, where $\theta$ is obtained by calculating the dot product of the previous timestep’s human action $a^{h}$ and shared action $a^{s}$.
+This assesses alignment, increasing $\gamma$ for positive alignment to enhance agent control, and reducing it for misalignment to increase user control.
+After obtaining the control ratio $\gamma$, we calculate the shared action $a^{s}$, using the human operator’s action $a^{h}$ as input, as defined in Eq. 8 (shown in Fig. 8).
+IV EXPERIMENTS
+In this section, we introduce the settings for both real world tasks and simulation tasks, along with the experimental results and data validation.
+Due to page limitations, we have included some detailed information such as training details and ablation study results on our webpage.
+IV-A Tasks.
 We adopt six multi-stage manipulation tasks (Fig.
-8).
+7).
 Pick-and-Place aims at picking an object on the table and placing it into a container.
@@
 Tool-Use aims at picking a hammer and using it to drive a nail into a board.
-Efficiency of Data Collection.
-Our proposed system leverages shared control between human operators and learned agents to enhance the efficiency of data collection.
-To learn how the assistant agent could improve the data collection process, we conducted a user study.
-Figure 8: Overview of six task settings and their task flow for Pick-and-Place (left), Articulated-Manipulation (middle), Gripper-Push (upper-right) and Dexterous-Tool-Use (bottom-right).
-In the user study, 10 human operators participate, collecting data under two modes: one where control is shared between the operator and the learned agent (w/ Ours), and the other where control is directly by the operator alone (w/o Ours).
-Each participant is instructed to collect as much data as possible within 3 minutes under two different modes for 3 dexterous hand tasks.
-Three metrics are evaluated: Success Rate (Percent) indicates the percentage of attempts where data collection was successful.
-Horizon Length (Steps per Sample) measures the length of each collected trajectory, with a lower horizon length indicating smoother data collection.
-Collection Speed (Samples per Hour) refers to the number of successful trajectories that can be collected in one hour.
 Pick-and-Place Door-Open Tool-Use
@@
 w/o Ours 45.42 471.48 120 53.45 317.21 176 34.47 511.03 124
-Table 1: User studies on three dexterous hand tasks.
+TABLE I: User studies on three dexterous hand tasks.
 Figure 11: Shared control process overview. The white one is the hand controlled purely by the human operator, while the cyan one is under shared control between the human and the assistive agent.
+IV-B Efficiency of Data Collection.
+Our proposed system leverages shared control between human operators and learned agents to enhance the efficiency of data collection.
+To learn how the assistant agent could improve the data collection process, we conducted a user study.
+In the user study, 10 human operators participate, collecting data under two modes: one where control is shared between the operator and the learned agent (w/ Ours), and the other where control is directly by the operator alone (w/o Ours).
+Each participant is instructed to collect as much data as possible within three minutes under two different modes for three dexterous hand tasks.
+Three metrics are evaluated: Success Rate (Percent) indicates the percentage of attempts where data collection was successful.
+Horizon Length (Steps per Sample) measures the length of each collected trajectory, with a lower horizon length indicating smoother data collection.
+Collection Speed (Samples per Hour) refers to the number of successful trajectories that can be collected in one hour.
 In Tab.
-1, by sharing control between humans and learned agents, our system shows improvements in both success rate and collection speed, while the average horizon length of the collected trajectories is reduced.
+I, by sharing control between humans and learned agents, our system shows improvements in both success rate and collection speed, while the average horizon length of the collected trajectories is reduced.
 This suggests that our system enhances the efficiency of data collection by facilitating a process that is easier to succeed, faster, and more fluid in terms of trajectory smoothness.
 To ensure the fairness of the experiment wasn’t compromised, we equally divided the user group into two parts, Group 1 first collected data directly by themselves (w/o Ours) and then collected data with an assistive agent (w/ Ours), while the Group 2 reversed the order, first (w/ Ours) mode and then (w/o Ours) mode.
-To gain deeper insight into how the learned agent assists the human operator, we visualize several keyframes from the data collection process of three dexterous hand tasks under shared control mode with $\gamma=0.4$.
+IV-C Quantitative Evaluation.
+To gain deeper insight into how the learned agent assists the human operator, we visualize several keyframes from the data collection process of three dexterous hand tasks.
 From Fig.
-9, it is evident that human operators are not required to provide too precise control with the assistive agent facilitating shared control over the dexterous hand.
+11, it is evident that human operators are not required to provide too precise control with the assistive agent facilitating shared control over the dexterous hand.
 Instead, they only need to convey high-level intentions, such as the direction of hand movement or finger grasp motions.
@@
 As a result, less effort and attention are required, making the data collection easier to execute successfully and speeding it up.
-Quantitative Evaluation.
-When the learned agent shares control with users, the system effectively corrects imperfect human control signals to achieve specific tasks.
-Given the challenge of directly measuring the level of imperfection in user signals and the correction ability of our system, we simulate the human using a baseline agent trained with Behavior Cloning (BC).
-Additionally, we import noise to the agent’s control signal through the diffusion policy’s forward diffuse process.
+When the learned agent shares control with users, the system effectively corrects imperfect human control signals to accomplish specific tasks.
+Given the challenge of directly measuring the level of imperfection in user signals and the correction ability of our system, we simulate human input using a baseline agent trained with Behavior Cloning (BC) as a proxy for user control.
 In Fig.
-LABEL:fig:simulateduser, the x-axis represents the forward ratio, where a higher forward ratio corresponds to the lower quality of action provided by the simulated operator and a higher control ratio of the learned agent on the dexterous hand.
+LABEL:fig:simulateduser, additional data collected under our framework effectively contributes to training improvements.
 The graph illustrates that with limited data availability, the agent can assist the simulated operator more effectively.
-As the agent accesses and trains with more data, its ability to correct actions improves.
-The results show that our system gradually diminishes the demand for the operator’s attention and effort, thereby enhancing the efficiency of data collection.
-Furthermore, once sufficient data is collected and the assistive agent is trained, it can transition into full autonomy mode by setting $\gamma$ to 1.0 and denoising random actions from the Gaussian distribution noise.
+As the agent gains access to and trains on more data, its ability to correct actions improves.
+These results indicate that our system gradually reduces the demand for the operator’s attention and effort, thereby enhancing the overall efficiency of data collection process.
+Furthermore, once sufficient data is collected and the assistive agent is trained, it can transition into full autonomy mode by setting $\gamma=1$ and denoising actions from the noise sampled from Gaussian distributions.
 Across three different dexterous manipulation tasks, we can achieve success rates of 0.76, 0.78, and 0.89, indicating that the assistive agent can effectively transform into an automated dexterous manipulation agent.
-From our experiments, we have observed that the assistive agent significantly aids human operators in managing fine control at the low level, especially in scenarios where accurate observation by humans is challenging, complicating effective action control.
+From our experiments, we have observed that the assistive agent significantly aids human operators in managing fine control, especially in scenarios where accurate observation by humans is challenging.
 For instance, tasks such as grasping an egg or moving a hammer present visual challenges.
@@
 DP 0.51 0.60 0.42 0.67 0.83 0.82
-Table 2: Data quality on downstream tasks.
+TABLE II: Data quality on downstream tasks.
 Dexterous Tool-Use Gripper Push-cube
@@
 […]
-Table 3: Tool-Use and Push-cube task success rate under increasing data.
-Data Quality on Downstream Task.
+TABLE III: Agent performance under increasing data.
+IV-D Data Quality on Downstream Task.
 In this section, we illustrate that collecting data under shared control does not compromise the quality of the data.
@@
 In Tab.
-2, compared to directly collecting human demonstrations from the expert human operator, who can achieve success rates and efficiency comparable to those with agent assistance, the data collected by sharing control between the human and the assistive agent can achieve comparable or even surprisingly better results with BC and BC-RNN.
+II, compared to directly collecting human demonstrations from the expert human operator, who can achieve success rates and efficiency comparable to those with agent assistance, the data collected by sharing control between the human and the assistive agent can achieve comparable or even surprisingly better results with BC and BC-RNN.
 Their results are comparable with DP, possibly as DP can better fit the tasks, which is in line with [8].
 In Tab.
-3, we compare the effects of using different sets of data to train BC and DP.
-We can find that utilizing more data collected under the shared control mode leads to comparable performance on the tool-use task.
+III, we compare the effects of using different sets of data to train BC and DP.
+We can find that utilizing more data collected under the shared control mode leads to comparable performance on the tool-use and push-cube tasks.
 This verifies that the new data contributes significantly to policy learning and can achieve a similar effect compared to the data from human experts but at a much lower cost.
 These results indicate that the data collected under the proposed paradigm have sufficient quality and efficiency for downstream tasks.
+IV-E Real World Experiment and User Feedback.
+To better evaluate our system, we further conduct real-world experiments.
+Three tasks are adopted: Pick-and-Place, Articulated-Manipulation, and Push-cube in Fig.
+14.
+Following the same rules as Sec.
+IV-B, four human volunteers are invited to participate in the user study to collect data under two modes: one where control is shared between the human operator and the learned agent (w/ Ours), and the other where control is directly by the human operator alone (w/o Ours).
+Our proposed system achieves significant improvements in success rate and collection speed by sharing control between human operators and learned agents, as demonstrated in Tab.
+IV.
+Additionally, data gathered under our proposed joint learning shared control mode yield performance on the three tasks that are comparable to those pure human datasets using BC and DP, further substantiated by the results presented in Tab.
+V.
 Figure 14: Real world setting.
 1.
-Pick-and-Place: (a) Hardware setup.
-The robot gripper should pick the red pot up (b-c) and place it onto the black induction cooker (d).
+Pick-and-Place: use the gripper to pick the red pot up and place it onto the black induction cooker.
 2.
-Articulated-Manipulation: use a gripper to open the drawer.
+Articulated-Manipulation: use the gripper to open the drawer.
 3.
-Push-cube: use a gripper to push the cube across the line.
+Push-cube: use the gripper to push the cube across the black line.
 Success Rate $\uparrow$ Horizon Length $\downarrow$ Collection Speed $\uparrow$
@@
 w/o Ours 0.70 21.54 121
-Table 4: Real world parallel gripper Pick-and-Place task user study.
+TABLE IV: Real world gripper Pick-and-Place task user study.
 Pick-and-Place Articulated-Manipulation Push-cube
@@
 DP 11 / 20 12 / 20 16 / 20 12 / 20 15 / 20 13 / 20
-Table 5: Real world parallel gripper experiments of data quality.
-Real World Experiment.
-To better evaluate our system, we further conduct real-world experiments.
-Three tasks are adopted: Pick-and-Place, Articulated-Manipulation, and Push-cube in Fig.
-15.
-Following the same rules as Sec.
-4, four human volunteers are invited to participate in the user study to collect data under two modes: one where control is shared between the human operator and the learned agent (w/ Ours), and the other where control is directly by the human operator alone (w/o Ours).
-Our proposed system achieves significant improvements in success rate and collection speed by sharing control between human operators and learned agents, as demonstrated in Tab.
-4.
-Additionally, data gathered under our proposed joint learning shared control mode yield performance on the three tasks that are comparable to those pure human datasets using BC and DP, further substantiated by the results presented in Tab.
-5.
-5 Conclusion
+TABLE V: Real world gripper experiments of data quality.
+Satisfaction: $\alpha=0.769$
+1. It is fun to use.
+2. It works the way I want it to work.
+3. It is wonderful.
+4. It helps me be more effective.
+5. It is flexible.
+User-Friendly: $\alpha=0.852$
+6. It is simple to use.
+7. It is effortless.
+8. I can use it without written instructions.
+9. I do not notice any inconsistencies as I use it.
+TABLE VI: Subjective Measures.
+We have developed a questionnaire comprising shown in Tab. VI to capture various dimensions of user experience and ergonomics, and we invited 10 volunteers to rate our system based on their feedback.
+This questionnaire assesses ease of use and overall satisfaction.
+The reliability of our questionnaire is supported by strong Cronbach’s alpha values: $\alpha=0.769$ for the satisfaction section and $\alpha=0.852$ for the user-friendly section, indicating internal consistency.
+V CONCLUSION
 In this paper, we introduce a novel human-agent joint learning paradigm that enables simultaneous human demonstration collection and robot manipulation teaching.
@@
 Our system offers a more efficient and flexible solution for data collection and robot manipulation learning via teleoperation.
-Acknowledgments
-This work is supported in part by the National Natural Science Foundation of China under Grants 62306175.